# Self-trapping and skin solitons in two-dimensional non-Hermitian lattices

### v1

@@ III ANTAGONISM BETWEEN SKIN EFFECT AND SELF-TRAPPING
For higher values of the non-Hermiticity parameter $h$, and different locations of the single-channel excitation, our results are presented in Fig. 5. More specifically, for $h=0.4$, Kerr nonlinearity can no longer induce self-trapping for single-site excitation at the center of the lattice, as shown in Fig. 5(a). Instead, the NHSE dominates, forcing the wavefunction to delocalize toward the corner (A) of the lattice. This behavior persists even as the excitation amplitude $A$ increases, within a reasonable range of physically relevant values, as indicated by the consistently low values of the $IPR$. Nevertheless, as shown in Fig. 5(b), the competition between nonlinearity and coupling asymmetry leads to different propagation characteristics when the initial excitation is near the corner (A) of the lattice, i.e., $\psi_{n_{x},n_{y}}(z=0)=A\delta_{n_{x},23}\delta_{n_{y},23}$. Contrary to the case of Fig. 5(a), self-trapping becomes possible above an amplitude threshold. This behavior can be attributed to the fact that the initial excitation is indeed near corner (A), towards which the couplings are higher for both spatial directions $n_{x}$ and $n_{y}$. At this point, it is interesting to also consider the case of an initial excitation near corner (B), i.e., at $(n_{x},n_{y})=(23,3)$. There, delocalization of wavefunctions is promoted only along the spatial dimension $n_{y}$, since the wavefunction is already on the side of the lattice towards which the coupling along $n_{x}$ directions higher. As shown in Fig. 5(c), although self-trapping is possible, it requires considerably higher excitation amplitude. This represents an intermediate case between those discussed in Fig.  5(a) and Fig.  5(b).
@@
To conclude, these examples highlight the intricate interplay between Kerr nonlinearity, which induces self-trapping, and non-Hermiticity, which promotes wavefunction delocalization. Up to this point, it is clear that the antagonism between those two effects depends on three different factors: the non-Hermiticity parameter $h$, the location of the initial excitation $(n_{x_{0}},n_{y_{0}})$, and its amplitude $A$. A natural and physically relevant question arising from our discussion is how to quantify such amplitude thresholds required to induce self-trapping at a single lattice site, for given non-Hermiticity parameter $h$.
IV Amplitude thresholds for Self-Trapping
@@
In order to investigate this question systematically, we study the dynamics in a 2D NLHN lattice, for single-site excitation at each site $(n_{x_{0}},n_{y_{0}})$, gradually increasing the excitation amplitude $A$. For each value of $A$, we compute the output $IPR$ (at $z_{\text{max}}=5$) of the wavefunction, $|\psi_{n_{x},n_{y}}(z_{\text{max}})|$, similarly with the previous section. We define the threshold for self-trapping as the lowest value of $A$ for which $IPR>0.8$. Such a criterion is meaningful as it signifies near-perfect self-trapping at a single site, and although the threshold choice is arbitrary, it does not compromise the generality of our results since any similar criterion would lead to identical conclusions. Furthermore, to ensure that the high value of $IPR$ arises from self-trapping rather than exponential localization toward the corner (A) of the lattice, we impose an additional constraint: the shift of the mean position of the wavefunction at $z=z_{\text{max}}$, $r\equiv\sqrt{(\langle n_{x}\rangle-n_{x_{0}})^{2}+(\langle n_{y}\rangle-n_{y_{0}})^{2}}$, must satisfy $r<1$. For reference, we mention that in the Hermitian case ($h=0$) the two aforementioned constraints yield an amplitude threshold of $A=3.15$ for all lattice sites $(n_{x_{0}},n_{y_{0}})$.
Figure 6: Amplitude thresholds $A$ (colormap) required for single- channel self-trapping for all possible site position. We consider two values of non-Hermiticity parameter: (a) $h=0.2$ and (b) $h=0.4$. Black squares in (b) indicate lattice sites where single-channel self-trapping is impossible for excitation amplitudes up to $A=20$.
@@
At first, we follow the previously described procedure for a lattice with $h=0.2$, and present our results in Fig. 6(a) as a comprehensive map. As evident from this figure, self-trapping occurs at relatively low excitation amplitudes, around $A\sim 3$, for lattice sites adjacent to corner (A). However, the threshold increases to approximately $A\sim 4$ for lattice sites near corners (B) and rises dramatically to $A\sim 9.5$ for the bulk of the lattice. When the non-Hermiticity parameter is increased to $h=0.4$ (Fig. 6(b)), the previously observed asymmetry in amplitude thresholds becomes significantly more profound. While sites near corner (A) still require an excitation amplitude of $A\sim 3$ for self-trapping, the threshold exceeds $A=10$ for sites near corner (B) and, more interestingly, becomes impossible for the bulk of the lattice. For this latter observation, we examined excitation amplitudes up to $A=20$. At this point we note, that when asymmetric couplings only occurs along one transverse direction, i.e., $h_{x}=0.4$, $h_{y}=0$, then self-trapping becomes feasible for all lattice sites.
@@
In conclusion, we have examined how Kerr nonlinearity induces self-trapping in a single-site when the excitation amplitude $A$ exceeds a specific threshold, depending on the location of the initial condition, in a 2D NLHN lattice. This analysis naturally leads to the question of whether this lattice also supports soliton solutions, analogous to the skin solitons recently predicted theoretically [46] and observed experimentally [47] in 1D NLHN systems.
V TWO-DIMENSIONAL SKIN SOLITONS
Figure 7: Power-eigenvalue $P-\mu$ diagrams for different families of solitons (a) near corner (A) $(n_{x_{0}},n_{y_{0}})=(23,23)$ and (b) in the bulk, for a 2D NLHN lattice. Colored lines correspond to different values of the non-Hermiticity parameter $h$. (c) Power thresholds $P_{th}$ as a function of $h$, for solitons near corner (A) (orange line) and in the bulk (purple line). Red, blue and green dots correspond to the dots of (a)-(b) indicating the value of $h$.
In this section, we identify stationary soliton solutions within the 2D NLHN lattice. In particular, we consider solutions of Eq. (1) ($g=1$) of the form $\psi_{n_{x},n_{y}}(z)=\phi_{n_{x},n_{y}}e^{i\mu z}$, where by $\phi_{n_{x},n_{y}}$ and $\mu$ we denote the soliton profile and the soliton eigenvalue, respectively. Substituting this ansatz into Eq. (1) results in a system of nonlinear algebraic equations,
$\begin{split}e^{h}\phi_{n_{x}-1,n_{y}}+e^{-h}\phi_{n_{x}+1,n_{y}}+e^{h}\phi_{n_{x},n_{y}-1}\\ +e^{-h}\phi_{n_{x},n_{y}+1}+|\phi_{n_{x},n_{y}}|^{2}\phi_{n_{x},n_{y}}=\mu\phi_{n_{x},n_{y}}.\end{split}$ (8)
@@ V TWO-DIMENSIONAL SKIN SOLITONS
Using iterative numerical techniques [52, 53, 54, 58], we obtain lattice soliton solutions and thus we can relate their corresponding power ($P\equiv\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\phi_{n_{x},n_{y}}|^{2}$) to the soliton eigenvalue ($\mu$). Such power-eigenvalue diagrams are shown in Fig. 7 for various values of $h$. Specifically, we analyze solitons positioned near corner (A) (Fig. 7(a)) and at the center of the lattice (Fig. 7(b)). In both cases, it is evident that the solitons exhibit power thresholds. Additionally, for both soliton locations, we investigate the dependence of the power thresholds $P_{th}$ on the non-Hermiticity parameter $h$. As illustrated in Fig. 7(c), the power thresholds for solitons located near corner (A) and at the center of the lattice increase monotonically with $h$. Specifically, we confirm that this dependence is parabolic, in good approximation, i.e., $P_{th}\approx ah^{2}+bh+c$. We also identify soliton solutions positioned near corners (B) and (C) of the lattice, having power–eigenvalue diagrams very similar to the solitons at the center of the lattice. For a given value of $h$, the power thresholds for corners (B) and (C) differ by less than $10^{-2}$ from those of the solitons in the center. However, the eigenvalues at these thresholds differ more, reflecting variations in the spatial distributions of the soliton profiles.
Figure 8: Soliton profiles $\big{|}\phi_{n_{x},n_{y}}\big{|}$ for $h=0.4$. Each panel corresponds to a different soliton solution family: (a) near corner (A) $(n_{x_{0}},n_{y_{0}})=(23,23)$, (b) near corner (B) $(n_{x_{0}},n_{y_{0}})=(23,3)$, (c) near corner (C) $(n_{x_{0}},n_{y_{0}})=(3,3)$, and (d) the center of the lattice $(n_{x_{0}},n_{y_{0}})=(13,13)$
Furthermore, in Fig. 8 we demonstrate the soliton profiles with the lowest eigenvalue at high non-Hermiticity parameters for different lattice locations. In all cases, the solitons exhibit spatial asymmetry toward corner (A), where linear modes are localized, due to the existence of asymmetric couplings. We also confirmed that, as one would expect, the degree of asymmetry of the soliton solutions increases when $h$ is increased.
@@
Finally, we investigate the stability of the soliton solutions by computing the deviation parameter defined as $\delta\phi_{n_{x},n_{y}}(z)\equiv\max\left(\big{|}|\phi_{n_{x},n_{y}}(z)|-|\phi_{n_{x},n_{y}}(0)|\big{|}\right)$ at a large propagation distance $z_{\text{max}}=25$. This parameter quantifies the deviation of the evolved soliton profile from its initial distribution. For a non-Hermiticity parameter $h=0.2$, solitons at all lattice locations remain stable, exhibiting very small deviations, with $\delta\phi\sim 10^{-7}$. However, at a higher non-Hermiticity value of $h=0.4$, soliton stability varies significantly with lattice position: solitons near corner (A) remain highly stable, with extremely low deviations ($\sim 10^{-10}$), while deviations increase progressively when moving from corner (B) ($\sim 10^{-7}$) toward the lattice center ($\sim 10^{-4}$), reaching their largest value near corner (C) ($\sim 10^{-3}$).
VI Conclusions and discussion
In this work, we have investigated the generalized nonlinear Hatano-Nelson lattice in two spatial dimensions, emphasizing the intricate interplay between skin effect and self-trapping due to Kerr nonlinearity. While in the linear regime, wavefunctions propagate toward a lattice’s preferred corner due to the NHSE, the presence of nonlinearity significantly affects such a preferential direction. We have demonstrated that the single-channel self-trapping strongly depends on the amplitude of the initial excitation, its position, and the degree of coupling asymmetry described. Specifically, initial excitations near the corner where the linear modes are localized, exhibit self-trapping above relatively low amplitude thresholds. Conversely, deep within the lattice bulk, asymmetric couplings actively oppose self-trapping, leading to delocalization even at higher excitation amplitudes; for very strong non-Hermiticity, self-trapping in these bulk regions is practically impossible. Additionally, we have identified and characterized stationary two-dimensional soliton solutions in various lattice locations by computing their associated power-eigenvalue diagrams. These solutions, which generalize the concept of skin solitons in two dimensions, consistently display spatial asymmetry biased toward the lattice corner which is favored by asymmetric couplings. Interestingly, our study revealed that the soliton power thresholds increase parabolically with the non-Hermiticity parameter $h$. Furthermore, solitons positioned near the corner where linear skin modes are localized, remain stable over large propagation distances, whereas those located away from this corner show higher deviations. In conclusion, this study introduces the concept of skin solitons in higher dimensional lattices with asymmetric coupling, setting the stage for further exploration of soliton formation in topologically non-trivial non-Hermitian lattices, such as those exhibiting higher-order skin effects. We believe that our results may provide useful insights regarding transport in higher-dimensional non-Hermitian lattices with nonlinearity, that are highly relevant to the emerging field of non-Hermitian nonlinear optics.
Acknowledgements.

### v2

@@ Results
Our results regarding higher values of the non-Hermiticity parameter $h$, and different locations of the single-channel excitation, are presented in Fig. 5. More specifically, for $h=0.4$, Kerr nonlinearity can no longer induce self-trapping up to $z=5$ for single-site excitation at the center of the lattice, as shown in Fig. 5(a). Instead, the NHSE dominates, forcing the wavefunction to delocalize toward the corner (A) of the lattice. This behavior persists even as the excitation amplitude $A$ increases, within a reasonable range of physically relevant values, as indicated by the consistently low values of the $IPR(z=5)$. Nevertheless, as shown in Fig. 5(b), the competition between nonlinearity and coupling asymmetry leads to different propagation characteristics when the initial excitation is near the corner (A) of the lattice, i.e., $n_{x_{0}}=n_{y_{0}}=23$. Since the linear eigenmodes of the system are localized near this location, self-trapping to the initial excitation-site becomes possible above an amplitude threshold, contrary to the case of Fig. 5(a). Interestingly, the pronounced peaks of $IPR(z=5)$ for $A\lesssim 4$ correspond to cases, where, after the wavefunction gets delocalized from the site of initial excitation due to coupling asymmetry, it becomes trapped in lattice sites closer to the corner (A) ($n_{x_{0}}=n_{y_{0}}=25$) of the lattice, for finite propagation intervals. At this point, it is interesting to also consider the case of an initial excitation near corner (B), i.e., at $n_{x_{0}}=23$ and $n_{y_{0}}=3$. There, delocalization of the wavefunction is promoted only along the spatial dimension $n_{y}$, since the wavefunction is already on the side of the lattice towards which the coupling along the $n_{x}$ direction is higher. As shown in Fig. 5(c), although self-trapping is possible, it requires considerably higher excitation amplitude. This represents an intermediate case between those discussed in Fig. 5(a) and Fig. 5(b).
@@
To conclude, the aforementioned examples highlight the intricate interplay between Kerr nonlinearity, which induces self-trapping of an initially localized wavepacket, and non-Hermiticity, which promotes wavefunction delocalization. Up to this point, it is clear that the antagonism between those two effects depends on three different factors; the non-Hermiticity parameter $h$, the location of the initial excitation $(n_{x_{0}},n_{y_{0}})$, and its amplitude $A$. A natural and physically relevant question arising from our discussion is how to quantify such amplitude thresholds required to induce self-trapping at a single lattice site, for a given non-Hermiticity parameter $h$.
Figure 5: Comparison of the self-trapping tendency for different locations of excitation channel, under fixed $h=0.4$. Normalized amplitudes $|\psi_{n_{x},n_{y}}(z=5)|$ for various values of AA (top row) under single-site excitation (a) at the center, $n_{x_{0}}=n_{y_{0}}=13$; (b) near corner (A), $n_{x_{0}}=n_{y_{0}}=23$; and (c) near corner (B), $n_{x_{0}}=23$ and $n_{y_{0}}=3$. IPRIPR for $z=5$ is shown as a function of AA in bottom row for all cases. Red dots indicate the values of $IPR$ for the corresponding values of AA in top row.
Amplitude thresholds for self-trapping
@@
In order to investigate this systematically, we study the dynamics in a 2D NLHN lattice, for single-site excitation at each site $(n_{x_{0}},n_{y_{0}})$, gradually increasing the excitation amplitude $A$. For each value of $A$, we compute the output $IPR(z=5)$ of the wavefunction, $|\psi_{n_{x},n_{y}}(z=5)|$, similarly to the previous subsection. We define the threshold $A_{th}$ for self-trapping as the lowest value of $A$ for which $IPR(z=5)>0.8$. Such a criterion is meaningful as it signifies near-perfect self-trapping at a single site, and although the threshold choice is arbitrary, it does not compromise the generality of our results since any alternative criterion would lead to similar conclusions. Furthermore, to ensure that the high value of $IPR(z=5)$ arises from self-trapping to the site of initial excitation rather than localization to other sites toward the corner (A) of the lattice, we impose an additional constraint: the shift of the mean position of the wavefunction at $z=5$, $r\equiv\sqrt{(\langle n_{x}\rangle-n_{x_{0}})^{2}+(\langle n_{y}\rangle-n_{y_{0}})^{2}}$, must satisfy $r<1$. For reference, we mention that in the Hermitian case ($h=0$) the two aforementioned constraints yield an amplitude threshold of $A=3.15$.
@@
At first, we follow the previously described procedure for a lattice with $h=0.2$, and present our results in Fig. 6(a) as a comprehensive map. As evident from this figure, self-trapping occurs at relatively low excitation amplitudes, around $A_{th}\sim 3$, for lattice sites adjacent to corner (A). However, the threshold increases to approximately $A_{th}\sim 4$ for lattice sites near corners (B) and rises dramatically to $A_{th}\sim 9.5$ for the bulk of the lattice. This dependence of $A_{th}$ on the excitation site stems from the position-dependent tendency for delocalization in the 2D HN lattice. Initial single-channel excitations at sites adjacent to corner (A), where couplings are stronger in both transverse directions, exhibit minimal delocalization even in the linear regime and therefore require the lowest amplitudes for self-trapping when nonlinearity is present. Excitations at sites near corner (B), where couplings are stronger along only one direction, exhibit important diffraction only along one direction and thus require higher thresholds, while excitations at bulk sites or near corner (C), which rapidly delocalize in both directions, require even larger amplitudes for self-trapping.
When the non-Hermiticity parameter is increased to $h=0.4$ [Fig. 6(b)], the previously observed asymmetry in amplitude thresholds becomes significantly more profound. While sites near corner (A) still require an excitation amplitude of $A_{th}\sim 3$ for self-trapping, the threshold exceeds $A_{th}=10$ for sites near corner (B) and, more interestingly, becomes impossible for the bulk of the lattice. For this latter observation, we examined excitation amplitudes up to $A=20$. At this point we note, that when asymmetric couplings only occurs along one transverse direction, i.e., $h_{x}=0.4$, $h_{y}=0$, then self-trapping becomes feasible for all lattice sites, as presented in the SM.
@@
In conclusion, in this subsection we have examined how Kerr nonlinearity induces self-trapping in a single-site when the excitation amplitude $A$ exceeds a specific threshold, depending on the location of the initially localized excitation and the degree of non-Hermiticity, in a 2D NLHN lattice.
Figure 6: Amplitude thresholds $A_{th}$ required for single-site self-trapping, for all lattice sites (colormap). Two values of non-Hermiticity parameter are considered: (a) $h=0.2$ and (b) $h=0.4$. Gray squares in (b) indicate lattice sites where single-site self-trapping is impossible for excitation amplitudes up to $A=20$.
Self-focusing and defocusing dynamics
Up to this point, we have examined the propagation dynamics in 2D NLHN lattices specifically for an initially localized single-channel excitation, uncovering a rich competition between self-trapping and lattice’s non-Hermiticity. As previously mentioned, for this type of excitation, the amplitude dynamics remain identical irrespective of the nature (focusing or defocusing) of the Kerr nonlinearity. Therefore, it is intriguing to investigate the dynamics under wider excitations to qualitatively explore the differences arising to the interplay of coupling asymmetry and Kerr nonlinearity, between the focusing ($g=1$) and defocusing (g=1g=-1) cases.
Specifically, we consider an initial Gaussian excitation centered at the lattice site $(x_{c},y_{c})$:
$\psi_{n_{x},n_{y}}(z=0)=A\frac{e^{-[(n_{x}-x_{c})^{2}+(n_{y}-y_{c})^{2}]}}{2w^{2}},$ (12)
where $A$ is the amplitude and $w$ is the Gaussian width parameter. In Fig. 7, we present an illustrative comparison between the propagation dynamics in the focusing (top row) and defocusing (bottom row) Kerr-nonlinear regimes for a non-Hermiticity parameter $h=0.2$. The initial excitation parameters are set to $w=2$, $(x_{c},y_{c})=(13,13)$, with the amplitude $A$ chosen to yield a total optical power of $\mathcal{P}_{\text{HN}}(z=0)=20$ [Fig. 7(a)/(d)]. Additionally, in Fig. 8, we plot the evolution of the wavefunction’s mean positions, $\langle n_{x}\rangle=\langle n_{y}\rangle$, and their uncertainties, $\Delta n_{x}=\Delta n_{y}$, comparing the focusing and defocusing nonlinear-regime propagation dynamics, with the corresponding linear case.
In the focusing nonlinear case, the wavefunction initially undergoes self-focusing toward the lattice center, as shown in Fig. 7(b), also evidenced by the decrease in the position uncertainties [Fig. 8(b)] and directly observed. Subsequently, however, the wavefunction begins to spread [Fig. 7(c)], and the mean positions begin to rise, though remaining consistently lower than in the linear scenario [Fig. 8(a)]. Notably, after a certain propagation distance, the wavefunction separates into two distinct parts, similarly to what discussed in the first subsection for single-site excitations: one fraction remains localized near the initial excitation location, while another part propagates toward corner (A) of the lattice [Fig. 7(d)], as reflected in increased position uncertainties for $z\gtrsim 5$, compared to the linear case [Fig. 8(b)].
Figure 7: Propagation dynamics in a 2D HN lattice for the focusing (top row) and defocusing Kerr-nonlinear (bottom row) cases, with non-Hermiticity parameter $h=0.2$ and initial condition $\psi_{n_{x},n_{y}}(z=0)=A{e^{-[(n_{x}-x_{c})^{2}+(n_{y}-y_{c})^{2}]}}/{2w^{2}}$, with $x_{c}=y_{c}=13$, $w=2$ and AA such that $\mathcal{P}_{\text{HN}}(z=0)=20$. (a)–(d) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the focusing case at propagation distances (a) $z=0$, (b) $z=1.5$, (c) $z=3$ and (d) $z=6$. (e)-(h) Normalized complex amplitude $|\psi_{n_{x},n_{y}}(z)|$ for the defocusing case at the same propagation distances.
In contrast to the single-site excitation scenario, the Gaussian initial condition [Eq. (12)] exhibits significantly different dynamics under defocusing Kerr nonlinearity. As illustrated in Figures 7 and 8, the wavefunction shifts away from its initial position markedly faster compared to the linear or focusing nonlinear cases, propagating preferentially toward corner (A) of the lattice, as reflected to increased mean positions. Importantly, it simultaneously exhibits stronger diffraction compared to the linear regime, as indicated by its higher position uncertainties. Thus, although the wavefunction mean position shifts toward the lattice’s preferred corner, the presence of coupling asymmetry does not prevent a significant fraction of the wavefunction from spreading broadly across the lattice.
Here, we note that the dynamics remain qualitatively similar to what described above for Gaussian excitations centered at different $(x_{c},y_{c})$ across the lattice. It is also noteworthy that, as $w$ decreases, the differences between the focusing and defocusing cases become less pronounced, and for $w\lesssim 0.3$ the propagation dynamics of the two regimes are almost indistinguishable, converging to the single-site excitation scenario.
Figure 8: Evolution of mean positions and uncertainties in a 2D HN lattice with non-Hermiticity parameter $h=0.2$, for initial condition $\psi_{n_{x},n_{y}}(z=0)=A{e^{-[(n_{x}-x_{c})^{2}+(n_{y}-y_{c})^{2}]}}/{2w^{2}}$, with $x_{c}=y_{c}=13$, $w=2$ and AA such that $\mathcal{P}_{\text{HN}}(z=0)=20$. (a) Mean positions $\langle n_{x}\rangle$ and $\langle n_{y}\rangle$ and (b) position uncertainties $\Delta n_{x}$ and $\Delta n_{y}$ for the linear case (gray lines), focusing Kerr-nonlinear (red lines), and defocusing Kerr-nonlinear cases (blue lines).
Two-dimensional skin solitons
In previous subsections, we systematically discussed that Kerr nonlinearity induces self-trapping of the wavefunction to its initial position, for single-channel excitation, when the excitation amplitude $A$ exceeds a threshold, which depends on the location of the initial excitation and the degree of coupling asymmetry. This analysis naturally leads to the question of whether this lattice also supports soliton solutions, analogous to the skin solitons recently predicted theoretically [55] and observed experimentally [56] in 1D NLHN systems. In this section, we identify stationary soliton solutions within the 2D NLHN lattice, particularly for focusing Kerr-nonlinearity ($g=1$). In particular, we consider solutions of Eq. (1) of the form $\psi_{n_{x},n_{y}}(z)=\phi_{n_{x},n_{y}}e^{i\mu z}$, where by $\phi_{n_{x},n_{y}}$ and $\mu$ we denote the soliton profile and the soliton eigenvalue, respectively. Substituting this ansatz into Eq. (1) results in a system of nonlinear algebraic equations,
$\begin{split}e^{h}\phi_{n_{x}-1,n_{y}}+e^{-h}\phi_{n_{x}+1,n_{y}}+e^{h}\phi_{n_{x},n_{y}-1}+e^{-h}\phi_{n_{x},n_{y}+1}+|\phi_{n_{x},n_{y}}|^{2}\phi_{n_{x},n_{y}}=\mu\phi_{n_{x},n_{y}}.\end{split}$ (13)
@@
Using iterative numerical techniques [63, 64, 65, 70], we obtain lattice soliton solutions and thus we can relate their corresponding power ($P\equiv\sum_{n_{x}=1}^{N}\sum_{n_{y}=1}^{N}|\phi_{n_{x},n_{y}}|^{2}$) to the soliton eigenvalue ($\mu$). Such power-eigenvalue diagrams are shown in Fig. 9 for various values of $h$. Specifically, we analyze solitons positioned near corner (A) [Fig. 9(a)] and near the center of the lattice [Fig. 9(b)]. In both cases, it is evident that the solitons exhibit power thresholds. We have also identified soliton solutions located near corners (B) and (C) of the lattice, having power–eigenvalue diagrams very similar to the solitons located near the center of the lattice. For a given value of $h$, the power thresholds for solitons near corners (B) and (C) differ by less than $10^{-2}$ from those of the solitons near the center. However, the eigenvalues at these thresholds differ more, reflecting variations in the spatial distributions of the soliton profiles.
Figure 9: Power-eigenvalue $P-\mu$ diagrams for different families of solitons (a) near corner (A) and (b) in the bulk, for a 2D NLHN lattice. Colored lines correspond to different values of the non-Hermiticity parameter $h$. (c) Power thresholds $P_{th}$ as a function of $h$, for solitons near corner (A) (blue line) and in the bulk (green line). Black, blue and red dots correspond to the dots of (a)-(b) indicating the value of $h$.
@@
Additionally, for both soliton locations, we investigate the dependence of the power thresholds $P_{th}$ on the non-Hermiticity parameter $h$. As illustrated in Fig. 9(c), the power thresholds for solitons located near corner (A) and near the center of the lattice increase monotonically with $h$. Specifically, we confirm that this dependence is well approximated by a parabolic form, $P_{th}\approx ah^{2}+bh+c$. In Fig. 10(a), we present a pertinent example of a soliton profile $|\phi_{n_{x},n_{y}}|$ located near corner (A), for $h=0.4$. In this case, as well as for solitons at other locations in the lattice, the solutions exhibit spatial asymmetry toward corner (A) due to the asymmetric couplings. As expected, we further verified that the degree of asymmetry of the soliton solutions increases with $h$. Also, we investigated the stability of the soliton solutions by computing the deviation parameter defined as $\delta\phi_{n_{x},n_{y}}(z)\equiv\max\left(\big||\phi_{n_{x},n_{y}}(z)|-|\phi_{n_{x},n_{y}}(0)|\big|\right)$, at a large propagation distance $z=25$. This parameter quantifies the deviation of the evolved soliton profile from its initial distribution. For a non-Hermiticity parameter $h=0.2$, solitons at all lattice locations remain stable, exhibiting very small deviations, with $\delta\phi\sim 10^{-7}$. However, at a higher non-Hermiticity value of $h=0.4$, soliton stability varies significantly with lattice position: solitons near corner (A) remain highly stable, with extremely low deviations ($\sim 10^{-10}$), while deviations increase progressively when moving from corner (B) ($\sim 10^{-7}$) toward the lattice center ($\sim 10^{-4}$), reaching their largest value near corner (C) ($\sim 10^{-3}$).
Finally, we examined the formation of skin solitons in other 2D lattices with asymmetric couplings under focusing Kerr nonlinearity, beyond the prototypical 2D NLHN lattice. As shown in Fig. 10, skin solitons also exist in alternative non-Hermitian lattices composed of properly engineered sublattices, where the global coupling asymmetry is directed, for example, toward a line, the lattice center, or other targeted regions [33]. These geometries are schematically illustrated in the left panels of Fig. 10(b)–(d). For each geometry, the right panels of Fig. 10(b)–(d) display a representative skin soliton localized near the region toward which the couplings are stronger, where the spatial asymmetry is clearly evident. For simplicity, all examples use 2D HN sublattices with non-Hermiticity parameter $h=0.4$. We numerically confirmed that skin solitons in these complex lattices also exhibit power thresholds. Consequently, the two-dimensional geometry offers a broad range of possibilities for supporting distinct asymmetric skin soliton solutions in complex non-Hermitian lattices, extending well beyond the 2D NLHN lattice [57].
Figure 10: Two-dimensional skin solitons in different lattice geometries. (a) 2D NLHN lattice with $N\times N=25\times 25$ sites and non-Hermiticity parameter $h=0.4$.. (b)–(d) Alternative lattices with asymmetric couplings: the directions of strong (weak) couplings are indicated by thick (thin) colored arrows in the left panels. At all boundaries between sublattices, the intra-sublattice couplings are reciprocal and set to $c=1$ (gray arrows). The right panels show representative soliton profiles $\big|\phi_{n_{x},n_{y}}\big|$ for the corresponding geometries, calculated for lattices with $N\times N=24\times 24$ sites composed of 2D HN sublattices with $h=0.4$. In each right panel, an inset displays a cross-section along the red dashed line.
Discussion
In this work, we have investigated the propagation dynamics of a generalized two-dimensional nonlinear Hatano–Nelson lattice, focusing on the intricate antagonism between the skin effect and self-trapping induced by Kerr nonlinearity. In the linear regime, wavefunctions are driven toward a lattice’s preferred corner by the NHSE, whereas the presence of nonlinearity substantially suppresses this delocalization tendency. In particular, for single-site excitation, we have shown that the self-trapping behavior strongly depends on the amplitude of the initial excitation, its position, and the degree of coupling asymmetry. Specifically, excitations near the corner where the linear modes are localized exhibit self-trapping above relatively low amplitude thresholds. By contrast, in the lattice bulk, asymmetric couplings effectively oppose self-trapping, leading to delocalization even at higher excitation amplitudes; for very strong non-Hermiticity, self-trapping in these bulk regions becomes practically impossible. Furthermore, to highlight the differences in this antagonism between focusing and defocusing Kerr-nonlinear regimes, we examined the propagation dynamics of broader initial excitations, revealing a stronger tendency for delocalization in the defocusing case, associated with enhanced diffraction.
Additionally, we have identified and characterized two-dimensional stationary bright soliton solutions in various lattice locations by calculating their associated power–eigenvalue diagrams. These solutions, which extend the concept of skin solitons in two dimensions, consistently display spatial asymmetry biased toward the lattice corner favored by asymmetric couplings. Interestingly, our study revealed that the soliton power thresholds increase parabolically with the non-Hermiticity parameter $h$. Remarkably, we have shown that such skin soliton solutions also exist for a plethora of asymmetric lattices with more complicated geometries, apart from the Hatano–Nelson lattice. Consequently, our results set the stage for further exploration of skin soliton formation in non-Hermitian lattices [57].
Importantly, recent advances in non-Hermitian photonics have enabled the experimental realization of nonlinear non-Hermitian systems in synthetic-mesh lattices [38, 56]. In this platform, Kerr nonlinearity has been implemented through an opto-electronic feedforward loop, enabling the observation of skin solitons in 1D HN lattices [56]. Moreover, the same synthetic-mesh architecture has been employed to realize linear 2D non-Hermitian lattices, where both spatial dimensions are synthesized by mapping fast and slow time bins of coupled fiber loops onto a two-dimensional lattice [38]. Given these developments, we believe that extending such schemes to implement Kerr nonlinear effects in 2D non-Hermitian lattices is within reach of current experimental capabilities. In conclusion, our results may provide useful insights into transport and soliton formation in higher-dimensional nonlinear non-Hermitian lattices, which are highly relevant to the emerging field of nonlinear non-Hermitian optics and photonics.
Acknowledgements